OpenStudy (anonymous):

Select the equation for the circle with center at point (–1, –8) and a radius of √5.

OpenStudy (anonymous):

(x+1)^2 +(y+8)^2 =(sqrt(5))^2 x^2 + 2x +1 +y^2 +16y +64 = 5 or x^2 + y^2 +2x +16y +60 =0